OpenStudy (anonymous):

using the pythagorean theorem (a squared + b squared = c squared) what is the length of the hypotenuse with legs the length of 6 and 8

OpenStudy (anonymous):

let a = 6 let b = 8 c^2 = a^2 + b^2 c^2 = 64 + 36 c^2 = 100 c = 10 make sense?